OpenStudy (anonymous):

Parker earned $36 by selling 35 glasses of fresh-squeezed lemonade at his lemonade stand on a Saturday morning. He charged $1.00 for a small glass and $1.50 for a large glass

pooja195 (pooja195):

et, parker sold x large glasses so, he sold (35-x) small glasses so, x*1.5 + (35-x)*1 = 36 or, 1.5x -x = 36-35 or, x = 1/.5 = 2 so he sold 2 large glasses